### v1

Abstract
The category was introduced in [12] in order to provide a structural setting to the study of the many Gelfand-Tsetlin subalgebras appearing in the context of the diagrammatic Soergel category of Elias and Williamson [3]. The category have as objects, all that we called nil graded algebras associated to a triangular matrix and as morphisms, all the preserving degree homomorphisms of graded algebras between them. In this article we develop a series of isomorphism criteria on that we used later to find out relevant information on the Gelfand-Tsetlin subalgebras of the diagrammatic Soergel category.
Contents
1 Introduction
1.1 Motivation
The diagrammatic Soergel Category was defined in 2016 by Elias and Williamson [3] in order to provide a diagrammatic approach to the study of the category of Soergel bimodules, a well known incarnation of the Hecke category, [18]. Since then, the category have gained attention of both algebraists and algebraic geometrist, due to its intrinsic connection with both branches of study.
A particular interest is the study of the internal structure of the category itself, that is, the structure of the many Hom spaces and more particularly of the many End algebras. In this context, for example, the diagrammatic setting provided by Elias and Williamson have been crucial to find out and prove the cellularity of the Light leaves bases of Libedinsky [8],[3]. Another example of the power of this diagrammatic approach is the discovery due to Ryom-Hansen, of explicit Jucys-Murphy elements relative to the Libedinsky’s bases [19].
In [10], Libedinsky and Plaza, found out a hard connection between the category and what they called the Blob category, another diagrammatic category, inspired in the well known KLR-algebras of Khovanov, Lauda and Rouquier. The similar patterns that they observed in both categories motivate them to define the Blob vs Soergel conjecture, where roughly speaking, they declare that there is an explicit incarnation of the category (in type ), in terms of KLR-diagrams.
Some of the first evidences on the veracity of the aforementioned conjecture can be found in the work of Espinoza and Plaza [4], where explicit isomorphisms were built between the corresponding Gelfand-Tsetlin subalgebras, that is, the algebras generated by the Jucys-Murphy elements, coming on one side from Two-color Soergel Calculus and on the other of certain idempotent truncation of the KLR version of the blob algebra (see [17]). In [13], Plaza, Ryom-Hansen and the author, obtain isomorphisms in a higher level, this time between certain endomorphism algebras of in type and the corresponding idempotent truncation of blob algebras. In both cases the procedure implies the search of optimal presentations in terms of generators and relations for the algebras involved, therefore, in both [4] and [13], the reader will find, not only a clear proof of each isomorphism, but also deep information on the structure of the corresponding algebras.
Recently in [1], Bowman, Cox and Hazi have proved the Blob vs Soergel conjecture, they in fact, proved a more general result, obtaining as a consequence the equivalences predicted by Libedinsky and Plaza. The smart procedure used in [1], based in alcove geometry and paths configurations, allows the authors to cover all the cases in just one go, while in [4] and [13] only particular cases were covered. On the other hand the general approach developed in [1] do not implies the aforementioned internal information on the corresponding algebras obtained in [4] and [13] or any generalization of that. Therefore, we consider as an open problem the search of this internal information for all the cases no covered in [4] and [13], that is in principle, to find out optimal presentations for all the algebras related with the Blob vs Soergel conjecture.
In the last few years, the author have been developing certain strategies to address this problem. For example in [11], presentations for the Gelfand-Tsetlin subalgebras of a family of idempotent truncations of the generalized blob algebras were found, later in [12], working on the side of the category the author developed a series of techniques and procedures to obtain optimal presentation for the Gelfand-Tsetlin subalgebras of the many End algebras for any type (far beyond the initial objective: type ). Both [11] and [12] are natural generalizations of [4] and particularly they seem to imply a hard connection between the light leaves bases of Libedinsky and the graded cellular bases built by Ryom-Hansen and the author in [14], just as was previously observed in [4] and [13], in that case, with respect to the graded cellular bases developed by Plaza and Ryom-Hansen in [17].
The category was introduced in [12] in order to provide a structural setting to the study of the many Gelfand-Tsetlin subalgebras appearing in the context of the diagrammatic Soergel category The category have as objects, all that we called nil graded algebras associated to a triangular matrix and as morphisms, all the preserving degree homomorphisms of graded algebras between them. Naturally, the aforementioned Gelfand-Tsetlin subalgebras of are (the main) objects in our category, and all the knowledge of developed in [12] allowed us to find out the desired presentations for those algebras. On the other hand, the category is by itself an interesting object of study and the results obtained in [12] open the door to a deeper analysis on it.
Being more specific, the objects of the category are commutative graded algebras defined by a presentation with a finite set of generators and certain quadratic relations codified by a strictly triangular matrix. In [12] we define and make a first study of those algebras, obtaining a series of techniques and corresponding applications to the study of the many Gelfand-Tsetlin subalgebras of the category Since each objects of is totally determined by a codifying triangular matrix, it seem plausible to ask when two of those objects are isomorphic and which role play the corresponding codifying matrices in this aspect. To answer this question is the main motivation of this and other works we have in progress.
In section 2, we develop a series of isomorphism criteria on that partially answer the above question. In Theorem 2.2 we obtain an absolute (but not easy to apply) matrix criterion to determine if two objects are or not isomorphic, this is the key result that allowed us to find out different approaches to a complete classification of the objects of Applying conveniently theorem 2.2, we could describe completely the class of the algebra associated to the zero matrix obtaining a very clear and easy to apply criterium (Theorem 2.9) to determine if the algebra is isomorphic to or not, depending on the entries of the codifying matrix The case of the matrix although very particular, can be cleverly used to determine if other two objects of are or not isomorphic (see for example the proof of proposition 3.3). Another fundamental result of this article is Theorem 2.39, where we describe another absolute criterium (again, not always easy to apply) to determine if two objects are or not isomorphic. In the case of theorem 2.39, the criterion states that two objects and are isomorphic if and only if, there is a finite sequence of certain matrix operations (that we called elementary triangular operations), that progressively transform the matrix into the matrix
Our first examples are developed in subsection 2.2, where the smaller cases were studied. The results that we found there inspire all the analysis we did in this article. In particular, they open interesting questions for further works, for example, since we observe that for small matrices, we have a very small number of isomorphic classes in then is natural to ask:
- Q1:
Is the number of classes always finite?
Although we do not have an answer for this question yet, in subsection 2.4, we obtain a lower bound for the number of classes depending on the size of the codifying matrices. In theorem 2.23, we describe an explicit set of representatives of different classes, for each size We left Q1 open, together with the search of a complete set of representatives and explicit criteria to describe any isomorphic class such as we did in the case of the matrix
On the other hand, in section 3, we advocate our attention to apply the results developed in section 2, to the study of the Gelfand-Tsetlin subalgebras coming from the diagrammatic Soergel category In [12], for each element we a define the SLTM, and the corresponding nil graded algebra We have proved in [12], that the algebra corresponds to the Gelfand-Tsetlin subalgebra of the endomorphism algebra in the diagrammatic Soergel Category of Elias and Williamson, relative to the Libedynski’s light leaves basis and the set of Jucys-Murphy elements defined by Ryom-Hansen (See [3],[8] and [19]).
In [12] we show that for the case of a Coxeter system of type the algebras are invariant under commuting moves. That result naturally open the following two questions:
- Q2:
Is the algebra invariant under commuting moves in any other type?
- Q3:
Is the algebra invariant under other braid moves?
In section 3, we answer those questions applying the different techniques developed in section 2. In theorem 3.2, considering the geometric realization of a Coxeter system (assuming all the technical conditions described in [3]), we prove that the algebra is invariant under commuting moves, independently of the particular type of the system. Although we already prove this for type in [12], the proof is different in the general case. In theorem 3.2 we use the new techniques developed in this article. Finally in proposition 3.3 we answer We provide an explicit example and again, we use the techniques developed in section 2 to verify that the algebra is not always invariant under other braid moves.
1.2 Acknowledgment
This work was supported by FONDECYT de Postdoctorado 2024, N°3240046. ANID, Chile.
1.3 Notations and terminology
is a fixed field of characteristic different from
is the zero matrix.
We denote by the matrix given by
(1.3.1) -
A SLTM, is a strictly lower matrix, with entries in A SLTM, is a strictly lower matrix (of any size), with entries in We denote by , the set of all the SLTM, and by , the set of all the SLTM.
If is a SLTM, and then we denote by the matrix obtained from by erasing the rows and the columns of from the th to the th.
Given a SLTM, a SLTM, and a matrix we define the matrix by blocks, as follows:
1.4 Generalities
In this subsection we recall the main facts on the category introduced in [12].
Definition 1.1.
Given a SLTM, the nil graded algebra associated to is the commutative unitary algebra defined by generators and relations:
We denote this algebra by We define a grading on by for any
Note that definition 1.1 implies that the algebra is a nil graded algebra, that is, a graded algebra, where any homogeneous element with degree different from is a nilpotent element (see [12] for details). In particular, if then definition 1.1 implies that is generated by square nilpotent elements.
We denote by the family of all nil graded algebras defied by some SLTM, and the family of all nil graded algebras defined by some SLTM.
The following theorem was proved in [12]:
Theorem 1.2.
If is a SLTM, then the dimension of the algebra is and the set is a basis for
Definition 1.3.
Given two SLTM, a preserving degree homomorphism is any homomorphism of algebras such that for any If also is an isomorphism we say that is a preserving degree isomorphism.
Definition 1.4.
We define the category as the one whose set of objects is and whose morphisms are all the preserving degree homomorphisms between them. An isomorphism in is a preserving degree isomorphism between two members of
In the following we shall say that the algebras and are isomorphic if they are isomorphic as object of the category (that is, in the sense of definition 1.4).
In this article we are interested in to find out certain matrix criteria, to determine when two objects of are isomorphic. Given two SLTM, we denote by whenever the algebras are isomorphic. Relation is a equivalence relation on the set of all SLTM.
2 Isomorphism criteria
2.1 The fundamental equations
We shall assume that is a SLTM, is a SLTM, and are the generators defining and respectively. A preserving degree homomorphism defines a matrix by the equation
The following lemma provides a first matrix criterium to find out if a given matrix defines a homomorphism or not.
Lemma 2.1.
Given a matrix, then the assignment
defines a preserving degree homomorphisms of algebras if and only if
Proof.
The assignment defines a morphism in the category if and only if for each we have:
If we develop the left side of equation 2.1.3, we obtain:
Now by relation 1.4.1, we obtain
On the other hand, if we develop the expression on the right in equation 2.1.3, we obtain:
If we develop the right side of equation 2.1.4, we obtain
Again by relation 1.4.1, we obtain
or equivalently:
for any and any pair such that as desired. ∎
The following theorem is a direct consequence of lemma 2.1. In this case we shall assume that and are both SLTM.
Theorem 2.2.
Given a matrix, then the assignment
defines a preserving degree isomorphisms of algebras if and only if
and is an invertible matrix.
2.2 Small cases
Proposition 2.3.
There is only one class in under relation
Proof.
Let
Then the matrix
defines an isomorphism ∎
The case is a little bit more complicated. We divide the analysis in a series of lemmas:
Lemma 2.4.
There is at least two classes in under relation
Proof.
Note that the matrices
are not related by In fact if there was an invertible matrix satisfying equation 2.1.6, then we have for any
If we assume that then by the second and third equations in 2.2.1, we have and but this contradicts the first equation in 2.2.1, therefore for all implying that is not invertible. ∎
Given a SLTM, we denote
Lemma 2.5.
Given a SLTM, such that and then we have
Proof.
1.
If then and the matrix
defines an isomorphism
2.
If then the matrix
defines an isomorphism
Lemma 2.6.
Given a SLTM, such that or then
Proof.
The following list shows all the different cases for satisfying the hypothesis of this lemma, and the corresponding matrix that defines an isomorphism
(Note that in the last case, we have since ). ∎
Proposition 2.7.
There are exactly two classes in under relation
Proof.
The results provided in this section invite us to ask ourselves two natural question:
- Q1:
Is the number of classes in (under relation ) always finite?
We do not have an answer for those questions yet, but with the techniques developed we are able to find at least a lower bound for that number, see subsection 2.4 below.
2.3 The class of
In this section we describe a very explicit criterium (theorem 2.9) for determining when a given SLTM, is in the same class of the matrix This criterium is a generalization of lemma 2.6.
We first need the following definition:
Definition 2.8.
Given a SLTM, a scalar and indices we define the triangular subdeterminant
Note that if we take for we recover the number defined in subsection 2.2.
Let and the generators of and respectively.
Theorem 2.9.
Given a SLTM, then if and only if one of the following conditions holds
1.
for each pair of indices
2.
Given a SLTM, then for each pair of indices such that we have that for each
Lemma 2.10.
If is a SLTM, such that then one of the following conditions holds
1.
for each pair of indices
2.
For each pair of indices such that we have that for each
Proof.
If is an isomorphism, then by theorem 2.2 we have
for each and each Lets assume that there is a for some pair with Since the matrix is invertible, there is a such that therefore equation 2.3.1 implies that
On the other hand if we take in equation 2.3.1, we can also conclude that therefore
since equation 2.3.1 is true for each for each and each Now equation 2.3.2 implies that
Therefore we have
Definition 2.11.
Let be a SLTM.
1.
A leader of is a pair such that and for all We denote by the set of all leaders of
2.
We define the graph of leaders of as the directed graph where the set of vertices is and where arrows are defined by:
We also denote by if there is a finite sequence such that and for each
3.
Let a connected subgraph of We denote by the set given by:
If the set is not empty, we say that it is a chain of leaders of If is a (not trivial) connected component of then we say that is a maximal chain of leaders. If are two different connected components of then we say that and are two independent chains of leaders of
4.
A simple chain of leaders of is a subset say with and such that
5.
A ramified chain of leaders of is any chain of leaders that is not simple.
Lemma 2.12.
Given a SLTM, we have that:
1.
Relation satisfy the following properties:
(2.3.3) and
(2.3.4) -
2.
The following relation:
(2.3.5) is a partial order on the set of vertices of
3.
For each there is a unique minimal element of (relative to ), such that Moreover if are in the same connected component of then there is a unique minimal element of such that and
4.
If in not a minimal element of under then the set
is a simple chain of leaders of
5.
is a minimal element if and only if there is no such that and that means that the th row of is zero.
The proof of lemma 2.12 is direct, we left to the readers.
Example 2.13.
1.
Let
(2.3.6) then In this case is itself a ramified chain of leaders of The associated graph looks like this:
Note that is composed by two simple chains: and
2.
Let
(2.3.7) Then In this case the associated graph is disconnected and it looks like this:
The sets are two independent simple chains of leaders such that
3.
Let
(2.3.8) In this case then is itself a simple chain of leaders. The associated graph in this case looks like this
4.
Consider the matrix
(2.3.9) In this case is itself a ramified chain, composed by two simple chains where
The associated graph in this case looks like this:
5.
Consider the matrix:
(2.3.10) Then the set is a ramified chain of leaders of The directed graph associated is given by:
Note that if we write and then are two (not independent) simple chain of leaders of such that
6.
Consider the matrix:
(2.3.11) Then the set is composed by two independent chains of leaders. The directed graph associated is given by:
Note that is divided in two independent (maximal) chains:
Also note that is a simple chain, while is a ramified chain, with where
Lemma 2.14.
Let be a SLTM. If for each pair with we have that for each then for each pair with such that we have that for each
Proof.
Let and suppose that In particular this implies that there is a such that By our hypothesis, for each and this implies that if and only if In particular, we have that This implies that in In fact, if we have that Otherwise, if then since there is a such that and again this implies that By a recursive argument, we can find a finite sequence: such that for each that is in other words
1.
Fix an If we assume that in then we have that:
The last equations imply that:
Therefore we have:
2.
Fix an . If there is a finite sequence: such that for each that is, then we have:
(2.3.12) (2.3.13) (2.3.14) From the equations in 2.3.12 we obtain respectively:
(2.3.15) Using recursively equations 2.3.13 and 2.3.14 we obtain respectively:
(2.3.18) and
(2.3.19) Finally we obtain:
as desired.
Lemma 2.15.
Assuming that is itself a simple chain of leaders, say such that, for each we have for each Then
Proof.
First note that the hypothesis implies that the presentation of is given by the relations
also the hypothesis implies that the all the entries of are determined by the values of with Since is a chain we have that for each We also denote
We define for each minimal element (in particular for ), the element Being minimal means that there is no such that in particular by lemma 2.12, we know that
For we define the element recursively by:
We claim that the assignment defines an isomorphism By equation 2.3.20, it is enough to prove that for each
For note that and then
For we assume that then
One can check recursively (similarly to the proof of lemma 2.14) that:
that is
and so on.
Therefore
and
Now note that condition for each implies that
Moreover, we can use equation recursively to check that
for any
In fact, it is clear that
Assuming that equation 2.3.25 is true, for any then we have:
therefore
and equation 2.3.25 is true for any for any
Example 2.16.
Let be the matrix of equation 2.3.8, then then the presentation of is given by the equations
Proposition 2.17.
Let be a SLTM, and assume that for each with we have that for each then
Proof.
For each minimal element we define Again, being minimal means that there is no a such that therefore lemma 2.12 implies that
For not minimal, we define recursively:
By property 2.3.4 of lemma 2.12, is well defined by equation 2.3.26.
By item 4 of lemma 2.12, we know that the set is a simple chain of leaders, therefore the fact that follows analogously to the proof of this in lemma 2.15. Note that each independent chain can be treated independently. Therefore we have that the assignment defines an homomorphism
It is clear that the associated matrix is invertible, therefore is an isomorphism. ∎
Example 2.18.
Example 2.19.
Example 2.20.
Let be the matrix of equation 2.3.9. As we mentioned in example 2.13 is a ramified chain, composed by two (not independent) simple chains, and One can check that satisfies all the conditions of proposition 2.17. The isomorphism described in the proof of proposition 2.17, is given by:
The associated matrix is
Example 2.21.
Let the matrix of equation 2.3.10. As we have mentioned in example 2.13, is a ramified chain composed by the two (maximal) simple chains: and Note that in this case the minimal vertices are all of them but are isolated vertices of Also note that satisfies the hypothesis of proposition 2.17. Therefore
Following the proof of proposition 2.17, we have that:
Using the recursive definition provided in equation 2.3.26 in the proof of proposition 2.17, we have for :
Analogously for we obtain:
2.4 A lower bound for the number of classes
In this subsection we show the following:
Theorem 2.23.
There are at least classes in .
To do so, we shall show that is a set of representatives of pairwise different classes under relation
Lemma 2.24.
Considering the matrix we have that if and only if
Proof.
Is not difficult to see that the matrix satisfies the conditions of theorem 2.9 if and only if ∎
Proposition 2.25.
The matrices and are related under if and only if
Proof.
By lemma 2.24, we know that whenever Therefore we only have to prove that the matrices and for are not related by
If we assume that there is an isomorphism then by theorem 2.2, we have
In particular if we fix a then for each pair equation 2.4.1 reduces to
If we assume that there is a such that then equation 2.4.2 implies that whenever and On the other hand, if we assume that for all then necessarily since But if we replace in equation 2.4.1, we obtain and therefore for all but this is a contradiction with the definition of the matrix The last analysis implies that for each there is a such that and for each with
Now we replace in equation 2.4.1, and such that then we obtain Since by definition of the matrix then we conclude that That is valid for each Therefore, since we conclude that By the last analysis we also conclude that for each pair we have In particular, we conclude that there are exactly rows such that
If we replace in equation 2.4.1 we obtain
and by definition of the matrices and the indices we obtain
Note that by the symmetry of equation 2.4.3 we can omit the hypothesis that
Going back again to equation 2.4.1 but this time taking we obtain
By definition of the matrix and by the facts that for each and we obtain more precisely that
If we take for some we obtain
Let us fix such that (we know there is at least two rows satisfying that condition. Then we have:
- 1.
- 2.
The last two cases implies that implies that and then there are at most rows such that but this is a contradiction, since we know there are exactly rows satisfying that condition. By the last analysis we conclude that and are not isomorphic, and then ∎
2.5 Elementary triangular operations
In this section we define a series of matrix operations called Elementary triangular operations (ETO), that preserves the class under relation At the end of the subsection we prove for any pair that for having is necessary and sufficient that there exists of a (finite) sequence of ETO that transform into
Definition 2.26.
Let be a SLTM, fix an and an scalar We define the matrix as the SLTM, where
Example 2.27.
Consider the matrix
then
Lemma 2.28.
Let ba a SLTM, fix and an scalar Then
Proof.
Let Let and the generators of and respectively. We claim that the assignment
defines an isomorphism To see that recall that by equation 2.1.3, we only have to check that and the associated matrix is invertible. Now note that
1.
If then clearly
2.
If then we have
and
3.
if then we have
and
By the last analysis, we conclude that the assignment defines an homomorphism Is not difficult to see that the associated matrix is invertible, in fact Therefore we conclude that defines an isomorphism ∎
Definition 2.29.
Let be a SLTM, and fix If the matrix satisfies the following conditions:
1.
The entry for each
2.
The entry for each
then we define the matrix as the SLTM, where
Example 2.30.
Consider the matrix
and take and Is not difficult to see that satisfies the conditions of definition 2.29. Then the matrix is well defined. Moreover we have
Lemma 2.31.
Let be a SLTM, let If satisfies the conditions of definition 2.29, then
Proof.
Let Let and the generators of and respectively. We claim that the assignment
defines an isomorphism To see that, note that by our hypothesis we have:
1.
For we have
and
2.
For we have
and
3.
For it is immediate that
4.
For we have
and
5.
For we have
and
The last analysis implies that the assignment defines a homomorphism Is not difficult to see that the associated matrix is invertible, in fact Therefore defines an isomorphism ∎
Recall that for a given SLTM, we denote
Definition 2.32.
Let be a SLTM, take an scalar and assume that there are indices such that for each If one of the following conditions holds:
1.
2.
and for each
Then we define the matrix as the SLTM, given by
Example 2.33.
Consider the matrix
Let and Note that the conditions of definition 2.32 hold. In this case Therefore we have
Example 2.34.
Therefore we have
Lemma 2.35.
Let a SLTM, let be a pair of indices and an scalar. Assuming that the conditions of definition 2.32 hold, then we have that
Proof.
Let Let and the generators of and respectively. We divide the analysis in two cases:
1.
If then we have that the matrix is given by:
We claim that the assignment:
defines an isomorphism In fact, if then it is clear that
We also can check for that
and
Finally for we have
and
Therefore, the assignment defines a morphism Is not difficult to see that the matrix is invertible, moreover then is an isomorphism.
2.
If and for each then the matrix as the SLTM, is given by
We claim that the assignment:
defines an isomorphism In fact, if then it is clear that
We also can check for that
On the other hand
Note that by our hypothesis, we have then
then as desired.
Finally for we have
and
Therefore, the assignment defines a morphism Is not difficult to see that the matrix is invertible, moreover then is an isomorphism.
Definition 2.36.
We call Elementary triangular operations (ETO for short) any of the following matrix assignments (whenever they are well defined):
Given two SLTM, and we say that they are equivalent by ETO if there is a finite sequence of ETO that transform into We denote by whenever and are equivalent by ETO.
Example 2.37.
Consider the matrices
then In fact one can check that
Definition 2.38.
In the following we shall denote respectively by the matrices given by:
Note that the matrix described in the proof of lemma 2.28 is of the form the matrix described in the proof of lemma 2.31 is of the form and the matrix described in the proof of lemma 2.35 is of the form
Theorem 2.39.
Let be two SLTM, then if and only if
Proof.
By lemmas 2.28, 2.31 and 2.35 we have that if then Moreover lemmas 2.28, 2.31 and 2.35 implies that for two SLTM, we have that if and only if there is an invertible matrix that can be written as a (finite) product of matrices of the form and
If then there exist an invertible matrix satisfying equation 2.1.6. Since is invertible, it can be factored as a product of elementary matrices (in the usual sense), Now the result follows from the fact that the matrices and are elementary matrices and any other elementary matrix can be written as a product of matrices of the form and ∎
3 Applications To Soergel Calculus
Fix a Coxeter system of finite rank (with for certain set of indices ), and a field where the number has sense in for each pair Following [6], the geometric realization of over a is a triple composed by a vector space subsets and where:
and the equations:
define representations of the group on and respectively. For any and any let
We are avoiding a technical definition on the diagrammatic Soergel category since it is well known. We invite the readers to look for more information in the original work of Elias and Williamson [3]. In this section we shall assume that the ground field has characteristic different from and satisfies all the technicalities described in [3]. Let the free monoid generated by the symbols In [12], for each element we define the SLTM:
and the corresponding nil graded algebra We have proved in [12], that the algebra corresponds to the Gelfand-Tsetlin subalgebra of the endomorphism algebra in the diagrammatic Soergel Category of Elias and Williamson (see [3]), relative to the Libedynski’s light leaves basis (see [8]) and the set of Jucys-Murphy elements defined by Ryom-Hansen in [19].
In [12] we show that for the case of a Coxeter system of type the algebras are invariant under commuting moves, and we proved that by defining explicit isomorphisms. Now in theorem 3.2 we generalize this fact for any other type. In this case we use the techniques developed in this article, to obtain a cleaner proof.
Let us recall first, certain technical definitions coming from [12]:
Definition 3.1.
1.
Let we define the column vector associated to as the matrix given by:
2.
Given two elements of say and we define the matrix by
where
In [12], we show that for any composed expression we have that:
Moreover in [12], we proved that for we have the following (kind of) associative property:
where
Now we are ready to prove the following:
Theorem 3.2.
For any Coxeter system the algebra is invariant under commuting moves.
Proof.
Let such that that its, the elements and of the group commute. Let and then are elements of that differ by a commuting move.
1.
Let and Then and Therefore:
(3.0.3) where
2.
Note that by definition and Therefore we have:
By the same reason, we have that two first rows of are given by:
analogously the two first rows of are given by:
3.
Note that for every we have that Let us denote
then
4.
Following with the notation of the s, we have that the rows of the matrices from the third to the last one are given respectively by:
and
But since and commute, we clearly have that
- 5.
Theorem 3.2 naturally opens the following question:
- Q3:
Is the algebra invariant under other braid moves?
Fortunately we have an answer for Q3, although it is not a positive one:
Proposition 3.3.
The algebra is not invariant under braid moves.
Proof.
It is enough to provide an example where two equivalent by braid moves expressions and generate two not isomorphic algebras and
Let be the Coxeter system of type Let and then:
1.
It is clear that and are equivalent by braid moves.
2.
The corresponding matrices and are given by:
3.
Considering the matrices defined by:
one can check that they are well defined (see subsection for details) and by definition More specifically we have:
4.
Let be the entries of let be the entries of the matrix By theorem 2.39, if the algebras and were isomorphic, then necessarily there will be an isomorphism Assuming that, then the associated matrix would satisfy equation 2.1.6 of theorem 2.2. In particular we would have for and any pair that:
(3.0.4) -
5.
Equation 3.0.4 in particular implies that:
- 6.
7.
Let and be the sets of generators of the algebras and respectively. From the last two items, we conclude that if there exists an isomorphism then necessarily for In particular this implies that the matrices and are in the same class under relation but this contradicts theorem 2.9. In fact
The analysis above implies that and are not isomorphic. ∎